sir kindly solve this qs density of monovalent sodium is 971kg/m3. find the density og charg carriers in it? at mass =0.023 kg/mol

sir
kindly solve this qs
density of monovalent sodium is 971kg/m3.
find the density og charg carriers in it?
at mass =0.023 kg/mol

Grade:12th pass

1 Answers

Soumya
5 years ago
for Na , M=23g/mol and ρ=0.971g/cm3
ne=NAρ/M= (6.02×1023electrons/mol)(0.971g/cm3) / 23g/mol
ne=2.54×1022 electrons/cm3= 2.54×1028 electrons/m3